# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ Appendix G Proof of Lemma 2
We extend the proof of the one-way coding architecture in [17] to the two-way coding architecture. Define the training data tuples as $\{\mathcal{T}_{j}\}_{j=1}^{J}$, where $\mathcal{T}_{j}=\{{\bf b}_{1}^{(j)},{\bf b}_{2}^{(j)},{\bf n}_{1}^{(j)},{\bf n}_{2}^{(j)}\}$ is the $j$-th tuple of the training data. Let us denote $\tilde{\eta}_{i}^{(j)}[k]$ as the output at timestep $k$ generated by data $j$, $\mathcal{T}_{j}$, through the User $i$’s encoding process of (59)-(61), $k=1,...,N$. It is obvious that $\tilde{\eta}_{i}^{(j)}[k]$ is independent and identically distributed (i.i.d.) over $j$ assuming the data tuples $\{\mathcal{T}_{j}\}_{j=1}^{J}$ are i.i.d. from each other. We define the mean and variance of $\tilde{\eta}_{i}^{(j)}[k]$ as $\mu_{i,k}$ and $\delta^{2}_{i,k}$, respectively. With the sample mean $m_{i,k}(J)=\frac{1}{J}\sum_{j=1}^{J}\tilde{\eta}_{i}^{(j)}[k]$ and the sample variance $d^{2}_{i,k}(J)=\frac{1}{J}\sum_{j=1}^{J}(\tilde{\eta}_{i}^{(j)}[k]-m_{i,k}(J))^{2}$, we define the normalization function as $\gamma_{i,k}^{(J)}(x)=(x-m_{i,k}(J))/d_{i,k}(J)$.
@@
Appendix H Derivation for $\Phi_{i}({\bf f}_{1,i})$ in (42)
We express each of the terms of $\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ in (35) as follows:
@@ Appendix H Derivation for $\Phi_{i}({\bf f}_{1,i})$ in (42)
With (84)-(87), we represent $\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ as
$\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ $=\sigma_{1}^{2}\bigg{(}\sum_{i=2}^{N-1}(q_{1,i-1}+f_{2,i}{\bf h}_{i}^{\top}{\bf f}_{1,i})^{2}+q_{1,N-1}^{2}+q_{1,N}^{2}\bigg{)}$
@@
$=({\bf h}_{1}^{\top}{\bf f}_{1,1})^{2}\sigma_{2}^{2}+\sum_{i=2}^{N-1}\bigg{(}\big{(}q_{1,i-1}+f_{2,i}{\bf h}_{i}^{\top}{\bf f}_{1,i}\big{)}^{2}\sigma_{1}^{2}+({\bf h}_{i}^{\top}{\bf f}_{1,i})^{2}\sigma_{2}^{2}+{\bf f}_{1,i}^{\top}{\bf f}_{1,i}(f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2})\bigg{)}$
$\hskip 28.45274pt+\sigma_{1}^{2}(q_{1,N-1}^{2}+q_{1,N}^{2}).$ (89)
@@
Appendix I Derivation of (44) from (43)
@@
Appendix J Derivation for $\alpha\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}+(1-\alpha)\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}$ in (50)
From the equation of $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (42), we can readily obtain
$\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}=2q_{1,i-1}{\bf h}_{i}^{\top}{\bf f}_{1,i}\sigma_{1}^{2}+2\sigma_{1}^{2}\big{(}|{\bf h}_{i}^{\top}{\bf f}_{1,i}|^{2}+\|{\bf f}_{1,i}\|^{2}\big{)}f_{2,i},\quad i\in\{2,...,N-1\}.$ (92)
For $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$, we first express each of the terms including ${\bf F}_{2}$ in $\mathbb{E}[\|{\bf x}_{2}\|^{2}]$ of (36) as a sum of entries of ${\bf F}_{2}$, i.e., $\{f_{2,i}\}_{i=2}^{N-1}$. First, revisiting the second term in (36), we obtain
$\|{\bf F}_{2}{\bf Q}_{1}^{1/2}{\bf q}_{1}\|^{2}$ $={\bf q}_{1}^{\top}{\bf Q}_{1}^{1/2}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf Q}_{1}^{1/2}{\bf q}_{1}={\bf p}^{\top}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf p}=\sum_{i=2}^{N}p_{i-1}^{2}f_{2,i}^{2},$ (93)
where we assumed that ${\bf p}\triangleq{\bf Q}_{1}^{1/2}{\bf q}_{1}=[p_{1},...,p_{N}]^{\top}$ is fixed for tractability although ${\bf Q}_{1}^{1/2}$ depends on ${\bf F}_{2}$. We then express the third term in (36) as
@@
$\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}=2p_{i-1}^{2}f_{2,i}+2\sigma_{1}^{2}f_{2,i}+2\sigma_{1}^{2}\bigg{(}\sum_{j=i+1}^{N-1}f_{1,j,i}^{2}f_{2,j+1}^{2}+\sum_{k=2}^{i-2}f_{1,i-1,k}^{2}f_{2,k}^{2}\bigg{)}+2\sigma_{2}^{2}\sum_{j=1}^{i-2}f_{1,i-1,j}^{2},$ (96)
for $i\in\{2,...,N-1\}$. Using (92) and (96), we obtain $\alpha\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}+(1-\alpha)\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}$ in (50).

### v2

@@ Appendix I Proof of Conjecture 1 When $N=3$
We first note that, if we choose ${\bf g}_{2}^{\star}=[0,...,0,\sqrt{\eta}\sigma_{2}]^{\top}$ (or equivalently, ${\bf q}_{2}^{\star}={\bf Q}_{2,\text{sqrt}}^{-1}{\bf g}_{2}^{\star}=[0,...,0,\sqrt{\eta_{2}}]^{\top}$), we have ${{\bf q}_{2}^{\star}}^{\top}{\bf B}{\bf q}_{2}^{\star}=(1-\alpha)\sigma_{2}^{2}\eta_{2}$ in (49) since all the entries in the last column and row of ${\bf B}$ are zeros except the last diagonal entry is $(1-\alpha)\sigma_{2}^{2}$. This choice of ${\bf q}_{2}$ may not be optimal when $\alpha<\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. When $\alpha<\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$, the lower bound in Conjecture 1 is $\alpha\sigma_{1}^{2}$. That is, we have $\alpha\sigma_{1}^{2}\leq\nu_{\min}[{\bf B}]$ for any ${\bf F}_{1}$ and ${\bf F}_{2}\in\mathcal{F}_{2}$. This means that the lower bound of the objective value ${\bf q}_{2}^{\top}{\bf B}{\bf q}_{2}$ in (49) (while satisfying the constraint $\|{\bf q}_{2}\|^{2}=\eta_{2}$) is $\alpha\sigma_{1}^{2}\eta_{2}$. Here, we can set the direction of ${\bf q}_{2}$ to align with the eigenvector corresponding to the smallest eigenvalue of ${\bf B}$. Therefore, in the worst case, i.e., when $\nu_{\min}[{\bf B}]=\alpha\sigma_{1}^{2}$ (with specific ${\bf F}_{1}$ and ${\bf F}_{2}$), the optimality loss in (49) will be $(1-\alpha)\sigma_{2}^{2}\eta_{2}-\alpha\sigma_{1}^{2}\eta_{2}=\eta_{2}(\sigma_{1}^{2}+\sigma_{2}^{2})(\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}-\alpha)$. ∎
Appendix L Proof of Lemma 3
Proof.
We extend the proof of the one-way coding architecture in [18] to the two-way coding architecture. Define the training data tuples as $\{\mathcal{T}_{j}\}_{j=1}^{J}$, where $\mathcal{T}_{j}=\{{\bf b}_{1}^{(j)},{\bf b}_{2}^{(j)},{\bf n}_{1}^{(j)},{\bf n}_{2}^{(j)}\}$ is the $j$-th tuple of the training data. Let us denote $\tilde{\eta}_{i}^{(j)}[k]$ as the output at timestep $k$ generated by data $j$, $\mathcal{T}_{j}$, through the User $i$’s encoding process of (80)-(82), $k=1,...,N$. It is obvious that $\tilde{\eta}_{i}^{(j)}[k]$ is independent and identically distributed (IID) over $j$ assuming the data tuples $\{\mathcal{T}_{j}\}_{j=1}^{J}$ are IID from each other. We define the mean and variance of $\tilde{\eta}_{i}^{(j)}[k]$ as $\mu_{i,k}$ and $\delta^{2}_{i,k}$, respectively. With the sample mean $m_{i,k}(J)=\frac{1}{J}\sum_{j=1}^{J}\tilde{\eta}_{i}^{(j)}[k]$ and the sample variance $d^{2}_{i,k}(J)=\frac{1}{J}\sum_{j=1}^{J}(\tilde{\eta}_{i}^{(j)}[k]-m_{i,k}(J))^{2}$, we define the normalization function as $\gamma_{i,k}^{(J)}(x)=(x-m_{i,k}(J))/d_{i,k}(J)$.
@@
Appendix M Derivation for $\Phi_{i}({\bf f}_{1,i})$ in (60)
We express each of the terms of $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}[\|{\bf x}_{1}\|^{2}]$ in (53) as follows:
@@
With (139)-(142), we represent $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}[\|{\bf x}_{1}\|^{2}]$ as
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}[\|{\bf x}_{1}\|^{2}]$ $=\sigma_{1}^{2}\bigg{(}\sum_{i=2}^{N-1}(q_{1,i-1}+f_{2,i}{\bf h}_{i}^{\top}{\bf f}_{1,i})^{2}+q_{1,N-1}^{2}+q_{1,N}^{2}\bigg{)}$
@@
$=({\bf h}_{1}^{\top}{\bf f}_{1,1})^{2}\sigma_{2}^{2}+{\bf f}_{1,1}^{\top}{\bf f}_{1,1}\sigma_{2}^{2}+\sum_{i=2}^{N-1}\bigg{(}\big{(}q_{1,i-1}+f_{2,i}{\bf h}_{i}^{\top}{\bf f}_{1,i}\big{)}^{2}\sigma_{1}^{2}+({\bf h}_{i}^{\top}{\bf f}_{1,i})^{2}\sigma_{2}^{2}$
$\hskip 28.45274pt+{\bf f}_{1,i}^{\top}{\bf f}_{1,i}(f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2})\bigg{)}+\sigma_{1}^{2}(q_{1,N-1}^{2}+q_{1,N}^{2}).$ (144)
@@
Appendix N Derivation of (62) from (61)
@@
Appendix O Derivation for $\alpha\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}+(1-\alpha)\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}$ in (68)
From the equation of $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (60), we can readily obtain
$\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}=2q_{1,i-1}{\bf h}_{i}^{\top}{\bf f}_{1,i}\sigma_{1}^{2}+2\sigma_{1}^{2}\big{(}|{\bf h}_{i}^{\top}{\bf f}_{1,i}|^{2}+\|{\bf f}_{1,i}\|^{2}\big{)}f_{2,i},\quad i\in\{2,...,N-1\}.$ (147)
For $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$, we first express each of the terms including ${\bf F}_{2}$ in $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}[\|{\bf x}_{2}\|^{2}]$ of (54) as a sum of entries of ${\bf F}_{2}$, i.e., $\{f_{2,i}\}_{i=2}^{N-1}$. First, revisiting the second term in (54), we obtain
$\|{\bf F}_{2}{\bf Q}_{1,\text{sqrt}}{\bf q}_{1}\|^{2}$ $={\bf q}_{1}^{\top}{\bf Q}_{1,\text{sqrt}}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf Q}_{1,\text{sqrt}}{\bf q}_{1}={\bf p}^{\top}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf p}=\sum_{i=2}^{N}p_{i-1}^{2}f_{2,i}^{2},$ (148)
where we assumed that ${\bf p}\triangleq{\bf Q}_{1,\text{sqrt}}{\bf q}_{1}=[p_{1},...,p_{N}]^{\top}$ is fixed for tractability although ${\bf Q}_{1,\text{sqrt}}$ depends on ${\bf F}_{2}$. We then express the third term in (54) as
@@
$\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}=2p_{i-1}^{2}f_{2,i}+2\sigma_{1}^{2}f_{2,i}+2\sigma_{1}^{2}\bigg{(}\sum_{j=i+1}^{N-1}f_{1,j,i}^{2}f_{2,j+1}^{2}+\sum_{k=2}^{i-2}f_{1,i-1,k}^{2}f_{2,k}^{2}\bigg{)}+2\sigma_{2}^{2}\sum_{j=1}^{i-2}f_{1,i-1,j}^{2},$ (151)
for $i\in\{2,...,N-1\}$. Using (147) and (151), we obtain $\alpha\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}+(1-\alpha)\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}$ in (68).
Appendix P Scatter plots of transmit symbols for linear and RNN-based coding
We compare the behavior of linear coding and RNN-based coding with scatter plots. We analyzed both coding schemes under different parameters: (i) $K=1$ and $N=3$ for linear coding in Fig. 17(17(a)) and for RNN-based coding in Fig. 17(17(b)), and (ii) $K=2$ and $N=6$ for linear coding in Fig. 18(18(a)) and for RNN-based coding in Fig. 18(18(b)). For RNN-based coding, we used the softmax function for decoding. We consider $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=10$dB. Each plot represents the scatter points of transmit symbols, ${\bf x}_{1}$ for User 1 and ${\bf x}_{2}$ for User 2, across the $N$ channel uses. For each bit-pair representation, we consider 1000 samples where each sample represents the randomness of noises ${\bf n}_{1}\sim\mathcal{CN}({\bf 0},10^{-\text{SNR}^{\text{ch}}_{1}/10}{\bf I})$ and ${\bf n}_{2}\sim\mathcal{CN}({\bf 0},10^{-\text{SNR}^{\text{ch}}_{2}/10}{\bf I})$, where we set the average power per channel use as $P=1$.
(a) Linear coding.
(b) RNN-based coding.
Figure 17: Scatter plots of transmit symbols across channel uses, when $K=1$, $N=3$, $\text{SNR}^{\text{ch}}_{1}=1$dB, and $\text{SNR}^{\text{ch}}_{2}=10$dB. The non-scattered point (the red point) indicates the transmit symbol used only for message transmission without containing any feedback information. It is interesting to see that some channel uses are not utilized in RNN-based coding, similar to linear coding. This may suggest that concentrating power on specific channel uses would be more effective for the scenario of $K=1$ and $N=3$.
We first explain the basic understanding of the scatter plots. In each plot, the non-scattered point (the red point) indicates no randomness at that channel use, meaning that the transmit symbol at that channel use does not contain any feedback information. Note that feedback information always contain the noise values, so it must be scattered in the scatter plot due to its randomness. Therefore, the non-scattered point represents message transmission only.
We first look at the case when $K=1$ and $N=3$ in Fig. 17. We consider that each bit $b_{i}\in\{0,1\}$ at User $i$ is modulated to $m_{i}=2b_{i}-1\in\{-1,1\}$ through pulse amplitude modulation (PAM). In Fig. 17(17(a)), we observe that User 1 uses the first and third channel uses, while User 2 uses the second and third channel uses. By Proposition 2, User 2 only transmits its message during the last channel use. User 1 transmits its message on the first channel use, receives feedback from User 2 on the second channel use, and then transmits a linear mixture of the message and feedback to User 2 on the third channel use. In Fig. 17(17(b)), for RNN-based coding, we observe that User 1 uses the first channel use to send its message and changes the sign of the transmit symbol based on the value of $b_{1}$: a negative value when $b_{1}=0$ and a positive value when $b_{1}=1$. Also, User 1 does not use the second channel use. These behaviors at User 1 are similar to those in linear coding. However, we observe that User 2’s behavior differs between linear and RNN-based coding. In RNN-based coding, User 2 uses the first channel use for its message transmission, while in linear coding, it uses the third channel use. Overall, it is interesting that some channel uses are not utilized in RNN-based coding, similar to linear coding. This may suggest that concentrating power on specific channel uses would be more effective for the scenario of $K=1$ and $N=3$
(a) Linear coding.
(b) RNN-based coding.
Figure 18: Scatter plots of transmit symbols across channel uses, when $K=2$, $N=6$, $\text{SNR}^{\text{ch}}_{1}=1$dB, and $\text{SNR}^{\text{ch}}_{2}=10$dB. For RNN-based coding in (b), the transmit symbols on the first three channel uses at User 1 are deterministic, and thus they contain only the transmit messages of ${\bf b}_{1}$ without any feedback information. User 2 uses the first and last channel uses to transmit its message of ${\bf b}_{2}$, which is a similar pattern to linear coding in (a). This pattern of the transmit symbols on the first and last channel uses at User 2 may be interpreted to be modulating two bits in ${\bf b}_{2}$ into transmit symbols.
Next, we look at the case when $K=2$ and $N=6$ Fig. 18. For linear coding, the alternate channel use strategy is employed. The first message at User $i$ is given by $m_{i}=2b_{i}[1]-1$, while the second message is $m_{i}^{\prime}=2b_{i}[2]-1$. For exchanging the first message pair $(m_{1},m_{2})$, User 1 uses odd-numbered channel uses, and User 2 uses even-numbered channel uses. For exchanging the second message pair $(m_{1}^{\prime},m_{2}^{\prime})$, User 1 uses even-numbered channel uses, and User 2 uses odd-numbered channel uses. In RNN-based coding in Fig. 18(18(b)), it is interesting to observe that the transmit symbols on the first three channel uses at User 1 are deterministic and do not exhibit any randomness. This indicates that these channel uses are only used for User 1 to transmit messages to User 2. It is also notable that User 2 uses the first and last channel uses to transmit its message, which is a similar pattern to linear coding. The signs of the transmit symbols on the first and last channel uses at User 2 are (+,-), (+,+), (-,-), (-,+) for each case of ${\bf b}_{2}=[0,0]$, $[0,1]$, $[1,0]$, $[1,1]$, respectively. Since RNN-based coding includes a modulation operation in its feedback coding, this pattern may be interpreted to be modulating two bits of ${\bf b}_{2}$ into transmit symbols.